OpenStudy (anonymous):

Factor completely: x2 - 49 (x - 7)(x - 7) (x + 7)(x - 7) (x + 7)(x + 7) Prime Isn't this one C?

OpenStudy (anonymous):

B..

jimthompson5910 (jim_thompson5910):

hint: a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

Thats what I meant sorry.

OpenStudy (anonymous):

I know how to multiply I swear. Lol

OpenStudy (camerondoherty):

lol

OpenStudy (anonymous):

use this a^2-b^2-(a+b)(a-b) here a=x and b=7

OpenStudy (camerondoherty):

@jim_thompson5910 already covered tht @zaibali.qasmi lol

OpenStudy (anonymous):

Differences of squares are factored using this formula. a^2−b^2=(a−b)(a+b) Since both terms are perfect squares, find the values for a and b. a^2=x^2, so a=x b^2=49, so b=7

OpenStudy (camerondoherty):

-,-

OpenStudy (anonymous):

(x-7)(x+7)